OpenStudy (anonymous):

Given the following curve, find (d^2 y/dx^2). 7x+2y=20

OpenStudy (anonymous):

first rewrite the equation where y is expressed as a function of x: y=-7x/2+10 then differentiate y'=-7/2 y''=0<answer

OpenStudy (anonymous):

out of my answer choices the best answer would be -49/4y^3. Right?

OpenStudy (anonymous):

what do you mean?